OpenStudy (anonymous):

assume that Assume a and b are nonzero rational numbers and c is an irrational number. For each of the following expressions, determine whether the result is irrational, rational, or both. Justify your answers. a(b + c) I think its irrational??

OpenStudy (dan815):

yep

OpenStudy (anonymous):

Ok I am wondering how to explain that?

OpenStudy (dan815):

rational + irrational = irrational and constants * irrational you can think of as adding irrational to itself multiple times for example 2*irrational = irrational+irrational 3*irrational = irrational + irrational + irrational

OpenStudy (dan815):

this fits into our first 3 rules we chose

OpenStudy (dan815):

irrational + irrational = irrational rational + irrational = irrational rational+rational = rational

OpenStudy (dan815):

just prove those 3 statements to yourself and then you can use those properties and time you want

OpenStudy (anonymous):

I see. thank you I was having a hard time understanding. Thank you so much

OpenStudy (anonymous):

So if we throw a radical in there how does it change?

OpenStudy (dan815):

also another useful thing you will need for your next question sqrt(prime) = irrational the square root of a prime number is irrational you can prove this by a contradiction suppose sqrt(prime) = rational sqrt(prime)=a/b, where a and b is the lowest fraction possible meaning gcd(a,b)=1 then prime = a^2/b^2 then b^2prime=a^2 but this means there are an odd number of factors on the left and and even number of prime factors on the right this is a contradiction, ergo the srt of a prime is not rational but irrational

OpenStudy (dan815):

some of thoses steps are not obvious at first sight, you might beed to investigate some of those steps further to make sense of it

OpenStudy (anonymous):

 So this too would be irrational?